can you explain where the t=sqrta comes from?can you explain where the t=sqrta comes from?
Answer #1

Given that " Thickness of Square beam of Douglas fir "

This indicates that beam is a square, so a=thickness * length = t*t =t^2 =Area

t^2 =A

so t=sqrt(A)